OpenStudy (anonymous):

How many times does the graph of y = x + 1 intersect the graph of y = x^2 + 3? Answer a. 1 b. 2 c. 3 d. 0 How many times does the graph of y = x + 1 intersect the graph of y = x^2 + 3? Answer a. 1 b. 2 c. 3 d. 0 @Mathematics

OpenStudy (anonymous):

I'm assuming y =–x2+3 means y=-x²+3 They are asking.. if the two equations equal each other how many solutions are there? -x²+3 = x+1 Put it all on one side: 0=x+1+x²-3 0=x²+x-2 x²+x-2=0 Solve the quadratic by factoring: (x-1)(x+2)=0 So x=1 or x=-2 Answer: Two times at x=1 and again at x=-2.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no Problem

OpenStudy (anonymous):

Wait, you shouldn't have three solutions... because that would imply that the function x + 1 - x^2 + 3 had three zeros which can't happen by the fundamental theorem of algebra

OpenStudy (anonymous):

sorry, x + 1 + x^2 - 3